# Across Platforms and Languages: Dutch Influencers and Legal Disclosures on Instagram, YouTube and TikTok

### v1

@@ 5 Findings
Legal Disclosure Practices. Overall, the amount of content voluntarily disclosed by influencers (green and yellow disclosures aggregated) shows that registered influencers only flag a marginal amount of their content as being monetized (5.63%) and, therefore, needing disclosure. Table 1 shows a general breakdown of the overall dataset and a distribution of disclosure practices and AM content across three platforms and two languages. Besides Dutch, the influencers also post content in English (43.5%).
@@
Percentage of disclosures 2.833% 3.393% 12.851% 7.298% 1.040% 2.518%
Green disclosure 1.039% 0.707% 0.009% 0.028% 0.135% 0.367%
Yellow disclosure 1.794% 2.686% 12.843% 7.270% 0.906% 2.151%
@@
Green disclosed AM 0.076% 0.050% 0.000% 0.001% 0.004% 0.004%
Yellow disclosed AM 0.164% 0.172% 1.294% 0.429% 0.046% 0.021%
Undisclosed AM 2.602% 0.238% 1.888% 12.346% 0.227% 0.115%
Within the disclosed content category, we note a very low usage of green disclosures in general, with YouTube having the lowest proportion (0.009%), where yellow disclosures are exclusively used (12.843%). One possible explanation is that green disclosures require strict positioning, so influencers may be placing them at the end of the text. On YouTube, this is additionally problematic since text on the platform tends to be longer than Instagram or TikTok posts. Overall, this finding reveals a preference of Dutch influencers for using popular disclosure cues that do not comply with Dutch law.
Table 1 illustrates the overall amount of AM content in the dataset per platform and language (total 4.76%), as well as how much AM is disclosed using green and yellow disclosures (total 0.43%). While green and yellow disclosures only allow us to track disclosures that were voluntarily made by influencers, they do not reveal non-disclosed advertising. Using the AM sub-dataset as a benchmark, it is possible to identify hidden advertising as non-disclosed AM (total 4.43%). Moreover, the green disclosure of AM content is meagre across all platforms and languages (even the highest is just 0.07% on Instagram English). Except for YouTube English, most AM content from the other venues remains undisclosed (especially for YouTube Dutch, with 12.346% of undisclosed AM content).
@@
% Green Disclosure 5.27 7.47 15.07 0.03 0.17 0.03 0.84 6.93 6.33
% Yellow Disclosure 17.33 31.00 23.86 4.53 79.16 16.09 10.75 35.84 39.31
% Green AM 1.06 0.69 2.12 0.00 0.01 0.00 0.00 1.01 1.01
% Yellow AM 2.01 4.24 4.13 0.06 3.84 4.11 0.00 7.07 9.09
% Non-disclosed AM 8.11 5.67 71.97 3.37 78.40 10.22 12.12 42.42 27.27
As suggested by Table 1, almost no green disclosures are found, and the majority (around 91%) of AM content stays undisclosed. These findings suggest that more disclosures originate from influencers with a large audience, whether in terms of overall disclosures or AM specifically. To further investigate this pattern, we analyze the top five influencers on each platform with the most AM disclosures. We then measure the proportion of disclosed AM among all AM created by each of the selected influencers, showing their compliance with disclosures. Fig. 2 presents the results.
@@
This plot suggests that for the micro-influencer group, no disclosure words are found positioned in the first five words of the sentence, which corresponds to the findings from Table 2. A general tendency for a higher median in disclosed AM content than in undisclosed ones can be observed except for micro-influencers on YouTube and mega-influencers on TikTok. Overall, these observations suggest that disclosures can benefit engagement but the results vary with the different positioning of disclosure words.
Lastly, we extend the experiment of the composition of different disclosures on Instagram from Fig. 1 and explore differences in engagement. “Green disclosures: word & position” are rarely found in all categories in Fig. 4 due to its few occurrences. Except for it and “green disclosures: toggle, words & position” in mega-influencers, green disclosures tend to perform better than yellow disclosures regarding the median. The variance of green disclosures is also better than other practices in micro- and mega-influencers. Moreover, in micro- and macro-influencers, “green disclosure: toggle, words & position” also performs better regarding the median than those only using toggle. The more compliant the AM posts are, the higher engagement they tend to attract. However, findings from mega-influencers contradict this assumption, as those using both toggle and words & position perform the worst.
@@ 6 Discussion and Future Research
This paper presents granular information on how disclosures are done on social media using Dutch law as a starting point to measure legal compliance. Our analysis shows that the general volume of disclosed content is astonishingly low. The content voluntarily disclosed by influencers, whether with green or yellow disclosures, amounts to a mere 5.63% out of the overall dataset. According to our results, in the case of affiliate marketing, only up to 9% is disclosed, leaving 91% of influencer marketing undisclosed. This result aligns with the low disclosure rates found in previous research on English YouTube and Pinterest affiliate marketing by [12], which was, on average, around 10%.

### v2

@@ 5 Findings
Legal Disclosure Practices. Overall, the amount of content voluntarily disclosed by influencers (green and yellow disclosures aggregated) shows that registered influencers only flag a marginal amount of their content as being monetized (6.53%) and, therefore, needing disclosure. Table 1 shows a general breakdown of the overall dataset and a distribution of disclosure practices and AM content across three platforms and two languages. Besides Dutch, the influencers also post content in English (43.5%).
@@
Percentage of disclosures 2.834% 3.708% 12.851% 10.146% 1.120% 3.013%
Green disclosure 1.049% 0.800% 0.011% 0.226% 0.143% 0.388%
Yellow disclosure 1.785% 2.908% 12.840% 9.920% 0.977% 2.624%
@@
Green disclosed AM 0.076% 0.055% 0.000% 0.027% 0.004% 0.004%
Yellow disclosed AM 0.164% 0.178% 1.294% 0.640% 0.046% 0.038%
Undisclosed AM 2.602% 0.227% 1.888% 12.110% 0.227% 0.098%
Within the disclosed content category, we note a very low usage of green disclosures in general, with YouTube English having the lowest proportion (0.011%), where yellow disclosures are exclusively used (12.840%). One possible explanation is that green disclosures require strict positioning, so influencers may be placing them at the end of the text. On YouTube, this is additionally problematic since text on the platform tends to be longer than Instagram or TikTok posts. Overall, this finding reveals a preference of Dutch influencers for using popular disclosure cues that do not comply with Dutch law.
Table 1 illustrates the overall amount of AM content in the dataset per platform and language (total 4.76%), as well as how much AM is disclosed using green and yellow disclosures (total 0.49%). While green and yellow disclosures only allow us to track disclosures that were voluntarily made by influencers, they do not reveal non-disclosed advertising. Using the AM sub-dataset as a benchmark, it is possible to identify hidden advertising as non-disclosed AM (total 4.27%). Moreover, the green disclosure of AM content is meagre across all platforms and languages (even the highest is just 0.076% on Instagram English). Except for YouTube English, most AM content from the other venues remains undisclosed (especially for YouTube Dutch, with 12.110% of undisclosed AM content).
@@
% Green Disclosure 5.51 7.62 14.90 0.52 0.71 0.09 0.72 5.97 6.17
% Yellow Disclosure 18.03 31.05 22.89 7.01 73.88 17.78 10.60 36.11 40.43
% Green AM 1.17 0.69 2.12 0.15 0.03 0.00 0.00 1.01 1.01
% Yellow AM 2.33 4.24 4.08 0.75 4.34 4.35 0.00 8.08 12.12
% Non-disclosed AM 7.68 5.67 72.02 2.52 77.87 9.97 12.12 41.41 24.24
As suggested by Table 1, almost no green disclosures are found, and the majority (around 90%) of AM content stays undisclosed. These findings suggest that more disclosures originate from influencers with a large audience, whether in terms of overall disclosures or AM specifically. To further investigate this pattern, we analyze the top five influencers on each platform with the most AM disclosures. We then measure the proportion of disclosed AM among all AM created by each of the selected influencers, showing their compliance with disclosures. Fig. 2 presents the results.
@@
This plot suggests that for TikTok few disclosures are found positioned in the first five words of the sentence for all sizes of influencers, which corresponds to the findings from Table 2. A higher median in late disclosure can be observed on Instagram, while early disclosure is higher than non disclosure for mega influencers on YouTube and TikTok. Overall, these observations suggest that disclosures can benefit engagement but the results vary with the different positioning of disclosure words.
Lastly, we extend the experiment of the composition of different disclosures on Instagram from Fig. 1 and explore differences in engagement. In Fig 4, “Green disclosures: word & position” are only visible in micro-influencers due to its few occurrences. Except for it and “Green disclosures: toggle, words & position” in mega-influencers, green disclosures tend to perform better than yellow disclosures regarding the median. The variance of green disclosures is also better than other practices in micro- and mega-influencers. Moreover, in micro- and macro-influencers, “Green disclosure: toggle, words & position” also performs better regarding the median than those only using toggle. The more compliant the AM posts are, the higher engagement they tend to attract. However, findings from mega-influencers contradict this assumption, as those using both toggle and words & position perform the worst.
@@ 6 Discussion and Future Research
This paper presents granular information on how disclosures are done on social media using Dutch law as a starting point to measure legal compliance. Our analysis shows that the general volume of disclosed content is astonishingly low. The content voluntarily disclosed by influencers, whether with green or yellow disclosures, amounts to a mere 6.53% out of the overall dataset. According to our results, in the case of affiliate marketing, only up to 10% is disclosed, leaving 90% of influencer marketing undisclosed. This result aligns with the low disclosure rates found in previous research on English YouTube and Pinterest affiliate marketing by [12], which was, on average, around 10%.